OpenStudy (anonymous):

A country's population in 1993 was 94 million. in 1999 in was 99 million. estimate the population in 2005 using the exponential growth formula. round your answer to the nearest millionth. P=Ae^kt

OpenStudy (cwrw238):

for the year 1999 P = 99 million A = 94 million , and t = 1999-1993 = 6 so we get 99000000 = 94000000 e^6k e^6k = 99/94 6k = ln (99/94) k = ( ln (99/94)) / 6 can you work that out with ur calculator ?

OpenStudy (cwrw238):

having found k you can find population in 2005 using P = 94,000000e^12k (t = 12 ( 2005 - 1993))